# Knowledge-aware Text-Image Retrieval for Remote Sensing Images

### v1

@@ I Introduction
However, the information asymmetry caused by cross-modality information is rarely mentioned, i.e. the fact that the short texts do not allow much freedom to represent diverse image content. When dealing with very high-resolution remote sensing images, the image content can be very diverse, hence it is difficult to be comprehensively summarized by a short caption only. On one hand, human captions can only describe one or a few aspects of the image, focusing on the most dominant information. For example, one image could receive the following caption text: There is a lake. Nevertheless, there might be trees and mountains around the lake which are ignored by humans or caption generators. On the other hand, different people will describe the image from subjective perspectives, resulting in a variety of text information for a single image, which may confuse the matching model. For example, the captions There is a lake and There are many boats could refer to the same image. Therefore, strategies to handle lacunary captions, nuances and synonyms are needed for the task, and a balance between objectivity and completeness must be achieved.
Commonsense knowledge sources have been recognized as effective priors in many vision-and-language research [13, 14] to reveal commonsense and alleviate ambiguities. For example, by adding commonsense knowledge from external sources, the description This is a lake will probably be extended by introducing related concepts and relations such as Lake has water and Boats on the lake, which extend the text content to match the image. In addition, external knowledge also provides an opportunity to link descriptions of different concepts. For example, the concepts lake and boat could be bridged by Boat on the lake from external knowledge sources.
@@ II-B External Knowledge Sources
Figure 2: (a) The pipeline of the KTIR. The proposed text-image retrieval system comprises three main components: an image encoder, a knowledge-aware text encoder and a similarity measurement module. The image feature ($\mathbf{f}_{img}$) are obtained by ViT [34] ($\operatorname{ViT}$). A BERT [35] is used as text-only mode ($\operatorname{BERT}_{text}$, green mode) and multimodal mode ($\operatorname{BERT}_{multi}$, yellow mode) to encode the knowledge-aware text feature ($\mathbf{f}_{txt}$) and the multimodal feature ($\mathbf{f}_{multi}$), respectively. Then the text-image contrastive loss ($\mathcal{L}_{\mathrm{con}}$) and the text-image matching loss ($\mathcal{L}_{\mathrm{mat}}$) are used as the training objectives for cross-modality retrieval. (b) The knowledge extraction process. In the knowledge-aware text encoder, the knowledge extraction process includes keyword extraction, knowledge retrieval and knowledge sentence construction. After knowledge extraction, the knowledge sentences $\mathbf{s_{k}}$ are collected for each caption $\mathbf{s}$.
@@ II-C Knowledge-aware Vision-Language Research
  • The knowledge sources: KCR only supports using RSKG as knowledge sources while KTIR also includes ConceptNet (See Section V-A).
@@ III-B1 Knowledge extraction
For a caption $\mathbf{s}$ with $n$ words: $\mathbf{s}=\{w_{1},w_{2},...,w_{n}\}$ $(n\geq 1)$, a tokenizer is used to separate every word and divide the part-of-speech (e.g. noun, verb, adjective, adverb, etc.). Based on the part-of-speech tags, all the nouns can be appended to a word list $\mathbf{s_{n}}$. Note that we turn all the plurals into their singular form.
@@
Detailed statistics of different external knowledge sources are shown in Table I. In general, $\mathbf{t_{rs}}$ has fewer objects and relation types compared to $\mathbf{t_{ce}}$ and $\mathbf{t_{co}}$. By combining two knowledge sources ($\mathbf{t_{co}}$), a reasonable amount of additional concepts and diverse types of relations are considered. Note that mining all the related objects and relations might be redundant, so in the experiments, we randomly select $m$ triplets from all the available ones for a caption $\mathbf{s}$ (Detailed analysis of triplet numbers can be found in Section V-D). In the end, the selected triplet list can be denoted as $\mathbf{t}$.
@@
Combine ($\mathbf{t_{co}}$) 908 41 748
Knowledge sentence construction
To encode the knowledge triplets, we use a template to convert them to knowledge sentences. More specifically, a knowledge triplet consists of a head, a relation and a tail, which are regarded as the subject, the predicate and the object of the short knowledge sentence, respectively. We try to rewrite the predicate to construct a meaningful sentence (e.g., UsedFor as “is used for”). In the end, a list of knowledge sentences is obtained based on the retrieved knowledge triplet list $\mathbf{t}$. Then we concatenate all knowledge sentences retrieved based on one caption together. The final knowledge sentence for caption $\mathbf{s}$ can be represented as $\mathbf{s_{k}}$.
@@ III-B2 Fusion in the text encoder
$\operatorname{CrossAtt}(\mathbf{f}_{1},\mathbf{f}_{2})=\operatorname{softmax}\left(\frac{W_{1}\mathbf{f}_{1}(W_{2}\mathbf{f}_{2})^{T}}{\sqrt{d_{k}}}\right)W_{2}\mathbf{f}_{2},$ (3)
where $d_{k}$ is the dimension of the feature space, and $W_{1}$ and $W_{2}$ are the transformation matrixes for the input features. We input both sequences (captions, then knowledge sentences and their reverse) as the input of the cross-attention layer. After being embedded in the feature space, a final FC layer is applied to obtain the overall representation of a text:
@@
where [ , ] denotes the concatenation operation.
@@
As stated in the previous subsection, the BLIP text encoder can also be used in multimodal mode ($\operatorname{BERT}_{multi}$), where the text features and image embeddings are used as the two inputs of the text encoder. The multimodal mode builds a joint feature ($\mathbf{f}_{multi}$) for multimodal inputs:
@@
Then the multimodal feature is sent to an FC layer ($\operatorname{FC}_{multi}$) to compute the probability of alignment between the image and text pair. This is implemented as a binary classification task: $p\left(y\right)=\operatorname{FC}_{multi}(\mathbf{f}_{multi})$, where $p\left(y\right)$ denotes the prediction of text-image pairs indicating whether they are matched or not.
@@ III-C Similarity Measurement
After encoding, the text and image are represented as image feature ($\mathbf{f}_{img}$), knowledge-aware text feature ($\mathbf{f}_{txt}$) and multi-modal joint embedding ($\mathbf{f}_{multi}$). In this section, we describe the objectives based on those features to perform cross-modal retrieval.
@@ III-C2 Knowledge-aware text-image matching loss
$\mathcal{L}_{\mathrm{mat}}=-\frac{1}{N}\sum_{i=1}^{N}y_{i}\cdot\log\left(p\left(y_{i}\right)\right)+\left(1-y_{i}\right)\cdot\log\left(1-p\left(y_{i}\right)\right),$ (8)
where $y_{i}$ is the binary label of image-text pair $i$ and $p\left(y_{i}\right)$ is the prediction of the model about whether the pair is a match. In practice, we follow the loss calculation method described in previous work [45, 30], using a momentum image encoder and a momentum text encoder with the same weights as those of the original encoders to produce features.
The final training objective $\mathcal{L}$ is the combination of knowledge-aware text-image contrastive loss and knowledge-aware text-image matching loss:
$\mathcal{L}=\mathcal{L}_{\mathrm{con}}+\mathcal{L}_{\mathrm{mat}}.$ (9)
@@ IV-A Datasets
  • UCM-Captions [15] is based on the UC Merced Land Use dataset [46]. It contains remote sensing images categorized into 21 scene categories (e.g., buildings, intersection, parking lost, runway, agricultural, forest, etc.), with 100 samples for each class. Each image contains 256 $\times$ 256 pixels and 5 sentences annotated.
  • RSICD [16] is a large remote sensing text-image dataset, and is a commonly used benchmark for remote sensing text-image retrieval and remote sensing image captioning. It contains 10921 images with the size 224$\times$224 pixels with various resolutions. There are also 5 sentences per image.
  • RSITMD [7] is a fine-grained remote sensing text-image retrieval dataset. It contains 4743 images from 24 categories with 23715 captions and 21403 keywords. Compared to the RSICD dataset, the RSITMD dataset was designed to have more fine-grained and diverse text descriptions.
@@ IV-C Implementation details
Following BLIP, the image encoder is a ViT-B/16, a ViT architecture with 12 attention heads, 12 hidden layers, and images divided into $16\times 16$ patches. The text encoder is $\text{BERT}_{\text{base}}$, a transformer encoder with 12 attention heads and 12 hidden layers. We initialize the encoder-decoder architecture with the corresponding pre-trained modules from BLIP [30]. Since all BLIP models are publicly available, we choose the “BLIP w/ ViT-B and CapFilt-L” checkpoint for initialization. This model was pre-trained on 129M noisy image-text pairs using CapFilt-L, a captioning and filtering method  [^2] https://github.com/salesforce/BLIP. In Section 9, we also integrate knowledge into CLIP [31]. For the CLIP-based model, we use ViT-B/32.
All the experiments were run on a single NVIDIA 4090. We train all the KTIR models for 10 epochs for each dataset, with a batch size of 16. We use AdamW as the optimizer and a cosine learning rate scheduler to adjust the learning rate during training. The starting learning rate is 5e-6, with a weight decay of 0.05. The temperature parameter ($\tau$) is initialized with 0.07. The CLIP-based models in Section 9 are trained for 30 epochs with a batch size of 64. We set the maximum number of triplets ($m$) as 5 for all the experiments, unless stated.
@@ IV-D Baseline Methods
We compare the proposed KTIR method with the following state-of-the-art methods in text-image retrieval. Based on the datasets that the model was trained on, there are two main categories: supervised training methods and pretraining-finetuning methods.
@@ IV-D1 Supervised training methods
  • KCR [44] proposes to use external knowledge triplets to extend the text content. It is the early version of KTIR. Note that, different from other methods, the text encoder is frozen in KCR.
@@ IV-D2 Pretraining-finetuning methods
V Results
V-A The Effectiveness of External Knowledge
In our first experiment, we investigate the effectiveness of the two different knowledge sources (RSKG and ConceptNet) and of their combination. In Table II, we compare KTIR with different knowledge sources with the fine-tuned BLIP modal on the UCM-Caption dataset. The results suggest external concepts can supply text information to fill the information gap between text descriptions and image content. For example, compared with the model without knowledge, KTIR with RSKG and ConceptNet improves 0.49% and 0.60% on mR, respectively. By combining the two sources, KTIR improves the performance of the baseline model by 1.14% on mR. Comparing the three different knowledge sources, KTIR with the combination of ConceptNet and RSKG outperforms the other two sources in isolation, suggesting the effectiveness of balancing external concepts and domain-specific knowledge.
Comparing the knowledge sentences by source
TABLE II: Experimental results on the UCM-Caption dataset when using different knowledge sources. Methods with external knowledge are reported with gray background. The best performances are marked in bold and the second best are underlined.
@@ V-A The Effectiveness of External Knowledge
A few examples of the retrieved knowledge sentences are shown in Fig. 4. Based on the keywords (in bold in the text), the knowledge retrieval process introduces relevant concepts (in green) to extend the text scope. From the figure we can see that both sources bring related objects or relations into knowledge sentences $\mathbf{s_{k}}$. For example, for the set of keywords (beach, sea, sands) in the first image, both sources add water to the knowledge sentences. Compared between those two sources, the knowledge sentences from the hand-crafted knowledge graph are more specific to the remote sensing field, while the information from the ConceptNet is more generic but diverse. For instance, in the second example, ConceptNet adds grass and garden that are highly related to the scene but it also includes computer that cannot be seen from the image. When combining two knowledge sources, the corresponding triplet can be filtered out as computer and house are not in RSKG.
@@
TABLE III: Results on the UCM-Caption dataset. Methods in the first block are the supervised learning methods, while in the second block, we present the pretraining-finetuning methods. * module weights frozen. Methods with external knowledge are reported with gray background. Note that we report the best results of comparing methods from the corresponding original papers on each dataset.
@@
TABLE IV: Rresults on the RSICD dataset. (See the caption of Table III for details)
@@
TABLE V: Results on the RSITMD dataset. (See the caption of Table III for details)

### v2

@@ I Introduction
However, the information asymmetry caused by cross-modality information is rarely mentioned, i.e., the fact that the short texts do not allow much freedom to represent diverse image content. When dealing with very high-resolution remote sensing images, the image content can be very diverse, hence it is difficult to be comprehensively summarized by a short caption only. On one hand, human captions can only describe one or a few aspects of the image, focusing on the most dominant information. For example, one image could receive the following caption text: There is a lake. Nevertheless, there might be trees and mountains around the lake which are ignored by humans or caption generators. On the other hand, different people will describe the image from subjective perspectives, resulting in a variety of text information for a single image, which may confuse the matching model. For example, the captions There is a lake and There are many boats could refer to the same image. Therefore, strategies to handle lacunary captions, nuances and synonyms are needed for the task, and a balance between objectivity and completeness must be achieved.
Commonsense knowledge sources have been recognized as effective priors in many vision-and-language research [13, 14] to reveal commonsense and alleviate ambiguities. For example, by adding commonsense knowledge from external sources, the description This is a lake will probably be expanded by introducing related concepts and relations such as Lake has water and Boats on the lake, which expand the text content to match the image. In addition, external knowledge also provides an opportunity to link descriptions of different concepts. For example, the concepts lake and boat could be bridged by Boat on the lake from external knowledge sources.
@@ II-B External Knowledge Sources
Figure 2: (a) The pipeline of the KTIR. The proposed text-image retrieval system comprises three main components: an image encoder, a knowledge-aware text encoder and a similarity measurement module. The image feature ($\mathbf{f}_{img}$) are obtained by ViT [34] ($\operatorname{ViT}$). A BERT [35] is used as text-only mode ($\operatorname{BERT}_{text}$, green mode) and multimodal mode ($\operatorname{BERT}_{multi}$, yellow mode) to encode the knowledge-aware text feature ($\mathbf{f}_{txt}$) and the multimodal feature ($\mathbf{f}_{multi}$), respectively. Then the text-image contrastive loss ($\mathcal{L}_{\mathrm{con}}$) and the text-image matching loss ($\mathcal{L}_{\mathrm{mat}}$) are used as the training objectives for cross-modality retrieval. (b) The knowledge extraction process. In the knowledge-aware text encoder, the knowledge extraction process includes keyword extraction, knowledge retrieval and knowledge sentence construction. After knowledge extraction, the knowledge sentences $\mathbf{s_{k}}$ are collected for each caption $\mathbf{s}$. Numbers in the feature vectors denote their dimension.
@@ II-C Knowledge-aware Vision-Language Research
  • The knowledge sources: KCR only supports using RSKG as knowledge sources while KTIR also includes ConceptNet (See Section 4).
@@ III-B1 Knowledge extraction
For a caption $\mathbf{s}$ with $n$ words: $\mathbf{s}=\{w_{1},w_{2},...,w_{n}\}$ $(n\geq 1)$, a tokenizer is used to separate every word and divide the part-of-speech (e.g., noun, verb, adjective, adverb, etc.). Based on the part-of-speech tags, all the nouns can be appended to a word list $\mathbf{s_{n}}$. Note that we turn all the plurals into their singular form.
@@
Detailed statistics of different external knowledge sources are shown in Table I. In general, $\mathbf{t_{rs}}$ has fewer objects and relation types compared to $\mathbf{t_{ce}}$ and $\mathbf{t_{co}}$. By combining two knowledge sources ($\mathbf{t_{co}}$), a reasonable amount of additional concepts and diverse types of relations are considered. Note that mining all the related objects and relations might be redundant, so in the experiments, we randomly select $m$ triplets from all the available ones for a caption $\mathbf{s}$ (Detailed analysis of triplet numbers and selection strategies can be found in Section V-D and Section V-E, respectively). In the end, the selected triplet list can be denoted as $\mathbf{t}$.
Knowledge sentence construction
Before encoding the knowledge triplets, we convert the selected ones (composing the list $\mathbf{t}$) to short knowledge sentences. More specifically, for a triplet ($<$head, relation, tail$>$), we keep the head subject and tail object as they are, but re-formulate the relation to construct a meaningful sentence according to a transformation template (e.g., UsedFor as “is used for”). The specific templates are listed in Table II. Following these rules, for example, the triplet $<$ boat, AtLocation, water $>$ can be re-written as boat is at location of water. After this re-formulation, a list of knowledge sentences is obtained, one per knowledge triplet in list $\mathbf{t}$. We then combine $m$ knowledge sentences as the final knowledge sentence for caption $\mathbf{s}$, which we refer to as $\mathbf{s_{k}}$.
@@
Combined ($\mathbf{t_{co}}$) 908 41 748
TABLE II: Templates used to convert relations in the knowledge triplets into knowledge sentences.
Relation Template Relation Template
UsedFor is used for shape shape is
ReceivesAction receives action color color is
HasA has a width width is
Causes causes distribution distribution is
HasProperty has a property height height is
CreatedBy is created by next_to next to
DefinedAs is defined as stop_at stop at
AtLocation is at location of pass_through pass through
HasSubEvent has intersect_at intersect at
MadeUpOf is made of marked_on is marked on
HasPrerequisite has prerequisite to connected_to is connected to
Desires desires is_component_of is component of
NotDesires not desires is_part_of is part of
IsA is a is_member_of is member of
CapableOf is capable of
@@ III-B2 Fusion in the text encoder
$\operatorname{CrossAtt}(\mathbf{f}_{1},\mathbf{f}_{2})=\operatorname{softmax}\left(\frac{\mathbf{f}_{1}W_{1}(\mathbf{f}_{2}W_{2})^{T}}{\sqrt{d_{k}}}\right)\mathbf{f}_{2}W_{2},$ (3)
where $d_{k}$ is the dimension of the feature space, and $W_{1}\in\mathbb{R}^{d_{k}\times d_{k}}$ and $W_{2}\in\mathbb{R}^{d_{k}\times d_{k}}$ are the transformation matrices for input features $\mathbf{f}_{1}\in\mathbb{R}^{d_{k}}$ and $\mathbf{f}_{2}\in\mathbb{R}^{d_{k}}$, respectively. We input both sequences (caption representations, then knowledge sentence representations and their reverse) as the input of the cross-attention layer. Note that after concatenation, the dimension of the concatenated sequences is reduced to $d_{k}$. After being embedded in the feature space, a final FC layer is applied to obtain the overall representation of a text:
@@
where [ , ] denotes the concatenation operation and the FC layer to reduce the feature dimensions.
@@
As stated in the previous subsection, the BLIP text encoder can also be used in multimodal mode ($\operatorname{BERT}_{multi}$), where the text inputs ($\mathbf{s,s_{k}}$) and image embeddings before $\operatorname{FC}_{img}$ layer, denoted as $\mathbf{f}_{img}^{\prime}$, are used as the two inputs of the text encoder. The multimodal mode builds a joint feature ($\mathbf{f}_{multi}$) for multimodal inputs:
@@
The multimodal feature is then used in a FC layer ($\operatorname{FC}_{multi}$) acting as a classifier to compute the probability of alignment between the image and text pair. This is implemented as a binary classification task:
$\hat{y}=\operatorname{FC}_{multi}(\mathbf{f}_{multi}),$ (6)
where $\hat{y}$ denotes the prediction for the text-image pair, indicating whether they are matched or not.
@@ III-C Similarity Measurement
After encoding, the text and image are represented as image feature ($\mathbf{f}_{img}$), knowledge-aware text feature ($\mathbf{f}_{txt}$), multi-modal joint embedding ($\mathbf{f}_{multi}$) and a predicted probability of matching ($\hat{y}$). In this section, we describe the objectives based on those features to perform cross-modal retrieval.
@@ III-C2 Knowledge-aware text-image matching loss
$\mathcal{L}_{\mathrm{mat}}=-\frac{1}{N}\sum_{i=1}^{N}y_{i}\cdot\log\left(\hat{y_{i}}\right)+\left(1-y_{i}\right)\cdot\log\left(1-\hat{y_{i}}\right),$ (9)
where $y_{i}$ is the binary label depicting whther the $i$-th image-text pair is a match and $\hat{y_{i}}$ is the corresponding predicted probability from the binary classifier (Eq. 6).
The final training objective $\mathcal{L}$ is the combination of knowledge-aware text-image contrastive loss and knowledge-aware text-image matching loss with the weights $w_{1}$ and $w_{2}$:
$\mathcal{L}=w_{1}\mathcal{L}_{\mathrm{con}}+w_{2}\mathcal{L}_{\mathrm{mat}}.$ (10)
In practice, we follow the loss calculation described in previous work [45, 30]: A momentum encoder is introduced to create soft labels as training targets. This is to account for the potential positives in the negative pairs for knowledge-aware text-image contrastive loss ($\mathcal{L}_{\mathrm{con}}$). We use a hard negative sampling strategy [45] to find negative samples showing the highest contrastive similarity in the mini-batch. We then use those samples for computing the knowledge-aware text-image matching loss ($\mathcal{L}_{\mathrm{mat}}$).
@@ IV-A Datasets
  • UCM-Captions [15] is based on the UC Merced Land Use dataset [46]. It contains remote sensing images categorized into 21 scene categories (e.g., buildings, intersection, parking lost, runway, agricultural, forest, etc), with 100 samples for each class. Each image contains 256 $\times$ 256 pixels and 5 sentences annotated.
  • RSICD [16] is a large remote sensing text-image dataset, and a commonly used benchmark for remote sensing text-image retrieval and remote sensing image captioning. It contains 10921 images with the size 224$\times$224 pixels with various resolutions. There are also 5 sentences per image.
  • RSITMD [7] is a fine-grained remote sensing text-image dataset. It contains 4743 images from 24 categories with 23715 captions and 21403 keywords. Compared to the RSICD dataset, the RSITMD dataset was designed to have more fine-grained and diverse text descriptions.
@@ IV-C Implementation details
Following BLIP, the image encoder is a ViT-B/16, a ViT architecture with 12 attention heads, $12$ hidden layers, and images divided into $16\times 16$ patches. The text encoder is $\text{BERT}_{\text{base}}$, a transformer encoder with $12$ attention heads and 12 hidden layers. The output feature dimension of both ViT and $\text{BERT}_{\text{base}}$ is $768$. The dimension of the final image feature ($\mathbf{f}_{img}$), of the knowledge-aware text feature ($\mathbf{f}_{txt}$) and of the multi-modal joint embedding ($\mathbf{f}_{multi}$) are $256$, $256$ and $768$, respectively. In the cross-attention mechanism, $d_{k}$ is set to $768$ as well, therefore matching the output of the text encoder. We initialize the encoder-decoder architecture with the corresponding pre-trained modules from BLIP [30]. Since all BLIP models are publicly available, we choose the “BLIP w/ ViT-B and CapFilt-L” checkpoint for initialization. This model was pre-trained on 129M noisy image-text pairs using CapFilt-L, a captioning and filtering method  [^2] https://github.com/salesforce/BLIP. In Section 9, we also integrate knowledge into CLIP [31]. For the CLIP-based model, we use ViT-B/32.
All the experiments were run on a single NVIDIA 4090. We train all the KTIR models for 10 epochs for each dataset, with a batch size of 16. We use AdamW as the optimizer and a cosine learning rate scheduler to adjust the learning rate during training. The starting learning rate is 5e-6, with a weight decay of 0.05. The temperature parameter ($\tau$) is initialized with 0.07 and is learnable during the training process. The CLIP-based models in Section 9 are trained for 30 epochs with a batch size of 64. We set the maximum number of triplets ($m$) as 5 for all the experiments, unless stated. The loss weights $w_{1}$ and $w_{2}$ are set to 1 experimentally. Their effect will be studied in Section V-D.
@@ IV-D Baseline Methods
We compare KTIR with the following state-of-the-art methods in text-image retrieval. Based on the datasets that the model was trained on, there are two categories: supervised training methods and pretraining-finetuning methods.
@@ IV-D1 Supervised training methods
  • KCR [44] proposes to use external knowledge triplets to expand the text content. It is the early version of KTIR. Note that, different from other methods, the text encoder is frozen in KCR.
@@ IV-D2 Pretraining-finetuning methods
TABLE III: Experimental results on the UCM-Caption dataset when using different knowledge sources. Methods with external knowledge are reported with gray background. The best performances are marked in bold and the second best are underlined.
@@
V Results
V-A The Effectiveness of External Knowledge
In our first experiment, we investigate the effectiveness of the two different knowledge sources (RSKG and ConceptNet) and of their combination. To further explore the effectiveness of different knowledge sources, we also augment KTIR using a large language model: we prompt GPT-4 [48] to generate relevant sentences to expand the scope of the given text, and then use these generated sentences as knowledge sentences for KTIR. To ensure a fair comparison, the number of generated sentences is kept the same as the number of knowledge sentences constructed from the other knowledge graphs, i.e., $m$. In Table III, we compare KTIR with different knowledge sources with the fine-tuned BLIP modal on the UCM-Caption dataset. The results suggest external concepts can supply text information to fill the information gap between text descriptions and image content. For example, KTIR with GPT-4, RSKG and ConceptNet improve the mR of the model without knowledge by 0.17%, 0.49% and 0.60%, respectively. By combining RSKG and ConceptNet, KTIR improves the performance of the baseline model even further, by 1.14%. The GPT-4-augmented KTIR underperforms KTIR with knowledge graphs. This is likely due to generative models hallucinating irrelevant sentences about the content, along with their lack of stability and reproducibility during the generation process, which might pose challenges to the retrieval task. Comparing the three different knowledge graphs as knowledge sources, KTIR with the combination of ConceptNet and RSKG outperforms the other two sources in isolation, suggesting the effectiveness of balancing external concepts and domain-specific knowledge. According to the experimental results, KTIR is not limited to a specific knowledge source and exhibits scalability for larger knowledge graphs. However, there is a trade-off between the size of the knowledge sources and the relevance of external knowledge. Especially, in the case of current remote sensing captioning datasets, the captions are generally short and span a limited set of concepts, which would advocate for smaller, more targeted knowledge graphs.
@@ V-A The Effectiveness of External Knowledge
Comparing the knowledge sentences by source
A few examples of the retrieved knowledge sentences are shown in Fig. 4. Based on the keywords (in bold in the text), the knowledge retrieval process introduces relevant concepts (in green) to expand the text scope. From the figure we can see that both sources bring related objects or relations into knowledge sentences $\mathbf{s_{k}}$. For example, for the set of keywords (beach, sea, sands) in the first image, both sources add water to the knowledge sentences. Compared between those two sources, the knowledge sentences from the hand-crafted knowledge graph are more specific to the remote sensing field, while the information from the ConceptNet is more generic but diverse. For instance, in the second example, ConceptNet adds grass and garden that are highly related to the scene but it also includes computer that cannot be seen from the image. When combining two knowledge sources, the corresponding triplet can be filtered out as computer and house are not in RSKG.
@@
TABLE IV: Results on the UCM-Caption dataset. Methods in the first block are the supervised learning methods, while in the second block, we present the pretraining-finetuning methods. * module weights frozen. Methods with external knowledge are reported with gray background. Note that we report the best results of comparing methods from the corresponding original papers on each dataset.
@@
TABLE V: Rresults on the RSICD dataset. (See the caption of Table IV for details)
@@
The impact of external knowledge
TABLE VI: Results on the RSITMD dataset. (See the caption of Table IV for details)